# CITRAS: Covariate-Informed Transformer for Time Series Forecasting

### v2

@@ 4.3 Model Analysis
4.3.1 Effect of Covariates
To illustrate the importance of covariates, we evaluate the performance of three top-ranking models, CITRAS, Timer-XL, and TimeXer, under three covariate usage settings: None (no covariate used), P (past covariates used), and P+F (past and future covariates used). Figure 3 demonstrates that relying solely on the target variable limits forecasting performance across all models, while covariates, particularly future covariates, play a crucial role in enhancing performance. Among these models, CITRAS exhibits the highest rate of improvement through the use of covariates, highlighting its effectiveness for covariate-informed forecasting. Visualization results that intuitively demonstrate the importance of covariates are in Appendix C.
Figure 3: Comparison of MSE under three variable usage settings: None (no covariate used), P (past covariates used), and P+F (past and future covariates used). The MSE change rates shown on the bar graph are calculated with respect to the "None" setting of each model.
4.3.2 Ablation Study
To elucidate the strong covariate understanding capability of CITRAS, we conduct the following ablation studies.
KV Shift
KV Shift captures the dependency between target variables (query) and known covariates (key) at the concurrent patch step $i$, and incorporates future known covariates (value) at step $i+1$. Another straightforward method to utilize future known covariates is to drop the first patch of known covariates and left-shift this channel entirely (denoted as Variate Shift), thereby evaluating the dependency between target variable at step $i$ (query) and future known covariates at step $i+1$ (key). However, Variate Shift imposes an additional burden on the model to reconcile the step misalignment in capturing the correlation, resulting in inferior performance compared to KV Shift, as shown in Table 3.
Table 3: Ablation Results of KV Shift. "Patch Step" represents the step of the key and value of the known covariate corresponding to the query of patch step $i$.
@@ 4.3.2 Ablation Study
0.320
Variate Shift
$i+1$
$i+1$
0.176
0.218
0.071
0.194
0.284
0.319
KV Shift
$i$
$i+1$
0.171
0.214
0.070
0.193
0.280
0.313
Attention Score Smoothing
Attention Score Smoothing captures locally accurate patch-level cross-variate dependency and refines it into global variate-level dependency. Note that a large alpha emphasizes patch-level dependency (with $\alpha=1.0$ corresponding to no smoothing), while a smaller alpha emphasizes variate-level dependency through heavier smoothing. Figure 4 (Left) illustrates the performance at different levels of granularity in cross-variate attention. The performance of the variate-level approach is represented by TimeXer. We observe that the suitability of variate-level versus patch-level approaches is dependent on the dataset. However, strong Attention Score Smoothing (around $\alpha=0.2$) strikes a balance between the two, highlighting the versatile effectiveness of this mechanism in both covariate-informed and multivariate forecasting.
4.3.3 Representation Analysis
To uncover the effects of Attention Score Smoothing further, we present the visualization of the cross-variate attention scores between the target variable and other variables in the EDF dataset in Figure 4 (Right). It is observed that patch-wise attention scores effectively capture the local negative impact of "Holiday" on the target variable ("Electricity Demand") . However, these scores are also susceptible to local disturbances, as demonstrated by the relatively fluctuating scores of "Humidity" compared to the "Temperature" that shows stable correlation with the target variable. By applying Attention Score Smoothing, the obtained scores effectively convey the strong dependency on "Holiday" from past steps and mitigate the noisy interaction caused by the fluctuating scores of "Humidity". In this way, Attention Score Smoothing enjoys the advantages of both patch-level and variate-level approaches while offering improved interpretability.
Figure 4: Left: The performance at different levels of granularity in cross-variate attention. The performance at "Variate" level is represented by TimeXer. Strong Attention Score Smoothing (around $\alpha=0.2$) outperforms both variate-level and patch-level approach. Right: Visualization of attention scores between the target variable ("Electricity Demand") and other variables observed in EDF dataset. By applying Attention Score Smoothing, the smoothened scores capture strong dependency on "Holiday" while mitigating noisy interaction between "Humidity".
@@ 5 Conclusion
Considering the critical roles of covariates in practical time series forecasting, we propose CITRAS, a decoder-only Transformer that flexibly leverages multiple targets, observed covariates, and known covariates. Specifically, we introduce two novel mechanisms within the patch-wise cross-variate attention modules: Key-Value (KV) Shift, which seamlessly integrates future values of known covariates into forecasting, and Attention Score Smoothing, which refines locally accurate dependencies into global variate-level dependencies. Experimentally, CITRAS outperforms state-of-the-art models on thirteen benchmarks from both covariate-informed and multivariate settings, demonstrating its versatile ability to leverage cross-variate and cross-time dependencies for improved forecasting accuracy. In the future, we will explore the generalization capability of CITRAS through large-scale pre-training to take advantage of its simple decoder-only Transformer architecture.
@@ A.1 Dataset Descriptions
We perform covariate-informed forecasting on seven real world datasets. EPF datasets [23] consist of five subsets from different day-ahead electricity markets, each spanning six years: (1) EPF-NP documents the Nord Pool electricity market from 2013-01-01 to 2018-12-24, containing hourly electricity prices as a target variable, with corresponding grid load and wind power forecasts as known covariates. (2) EPF-PJM records the Pennsylvania-New Jersey-Maryland market from 2013-01-01 to 2018-12-24, containing zonal electricity prices in the Commonwealth Edison (COMED) as a target variable, with corresponding system load and COMED load forecasts as known covariates. (3) EPF-BE captures Belgium’s electricity market from 2011-01-09 to 2016-12-31, containing hourly electricity prices as a target variable, with corresponding load forecasts in Belgium and generation forecasts in France as known covariates. (4) EPF-FR documents the electricity market in France from 2012-01-09 to 2017-12-31, containing hourly prices as a target variable, with corresponding load and generation forecasts as known covariates. (5) EPF-DE records the German electricity market from 2012-01-09 to 2017-12-31, containing hourly prices as a target variable, with zonal load forecasts in the TSO Amprion zone and wind and solar generation forecasts as known covariates. (6) EDF dataset [2] is an hourly time series dataset spanning over five years from 2015-01-03 to 2020-06-27, where the target variable is the national electricity demand in Panama. Observed covariates include four meteorological indicators (T2M=Air temperature at 2 m (${}^{\circ}\mathrm{C}$), QV2M=Specific humidity at 2 m (%), TQL=Total precipitable liquid water ($1/m^{2}$), and W2M=Wind speeed at 2 m ($m/s$)) measured in three cities (Tocumen, Santiago, David) in Panama, resulting in a total of 12 indices. In addition, we adopt three binary calendar features (holiday, school day, weekend) as known covariates. (7) BS [17] is an hourly time series data spanning two years from 2011-01-01 to 2012-12-31, where the target variables are three types of rental counts (casual, registered, and cnt=total count) in Bike Sharing system in Washington D.C. Observed covariates include five meteorological indicators (weathersit=weather situation, temp=temperature, atemp=feeling temperature, hum=humidity, and windspeed). In addition, we adopt three calendar features (holiday, weekday, working day) as known covariates.
@@
In addition, we perform multivariate forecasting on nine real-world datasets: ETT [56] contains 7 monitoring factors in the electricity transformers from July 2016 to July 2018. (1) ETTh1 and (2) ETTh2 are hourly subsets, and (3) ETTm1 and (4) ETTm2 are subsets recorded every 15 minutes. (5) Weather [49] includes 21 meteorological factors recorded every 10 minutes from the Weather Station of the Max Planck Biogeochemistry Institute in 2020. (6) ECL [49] comprises hourly electricity consumption data from 321 clients. (7) Traffic [49] records hourly road occupancy rates measured by 862 sensors on San Francisco Bay area freeways from January 2015 to December 2016. (8) PEMS04 [25] contains the public traffic network data recorded at 307 locations in California every 5 minutes from 2018-01-01 to 2018-02-28. Similarly, (9) PEMS08 [25] contains the data at 170 locations from 2016-07-01 to 2016-08-31. We use all the variables in these datasets as target variables. The summary of datasets are provided in Table 4.
@@
EPF-NP 1 0 2 1 Hour (36500, 5219, 10460)
EPF-PJM 1 0 2 1 Hour (36500, 5219, 10460)
Covariate-Informed EPF-BE 1 0 2 1 Hour (36500, 5219, 10460)
Forecasting EPF-FR 1 0 2 1 Hour (36500, 5219, 10460)
EPF-DE 1 0 2 1 Hour (36500, 5219, 10460)
@@
ETTh 7 0 0 1 Hour (8545, 2881, 2881)
@@
Multivariate ECL 321 0 0 1 Hour (18317, 2633, 5261)
Forecasting Weather 21 0 0 10 Minutes (36792, 5271, 10540)
@@ A.2 Experiment Settings
All the experiments are implemented in PyTorch [36]. We utilize a single NVIDIA A100 80GB GPU for the ECL, Traffic, and PEMS datasets and a single NVIDIA V100 32GB GPU for the other datasets. We apply Series Stationarization [31] to the embedding and projection layers of CITRAS to mitigate nonstationarity. We adopt ADAM [22] with an initial learning rate of $10^{-4}$ and L2 loss for model optimization. The training process is fixed at 10 epochs with an early stopping mechanism, which terminates training if the validation performance does not improve for 3 consecutive epochs. We set the number of layers in our proposed model $L\in\{1,2,3,4\}$, the number of heads in the multi-head attention layers as 8, dimension of token embedding $D\in\{128,256,512,1024\}$, and the smoothing factor for Attention Score Smoothing $\alpha\in\{0.1,0.2,0.4,0.8\}$. The patch length is uniformly set to 24 for covariate-informed forecasting and 96 for multivariate forecasting. We ensure that the drop last option is set to False in the test data loader to avoid any unfair impact, as highlighted in [37]. For the ablation study, we use the same hyperparameters as the full model to ensure a fair comparison.
@@ A.3 Baseline Descriptions
We reproduced the baseline models based on the TimesNet [48] repository or official repositories. Note that Timer-XL, TimeXer  [^1] The TimeXer implementation that utilizes future covariates is not officially released., iTransformer, and Leddam are extended from their original implementations to support future covariates in covariate-informed forecasting. Specifically, for Timer-XL, we modified the masking scheme in TimeAttention so that target token embeddings can attend to known covariate token embeddings from one patch step ahead. For TimeXer, iTransformer, and Leddam, we introduced an additional variate embedding layer for known covariates to handle their longer lengths.
@@ Appendix B Extra Model Analysis
B.1 Comparison between Decoder-only and Encoder-only Designs
CITRAS achieves the integration of covariates while preserving the canonical decoder-only Transformer architecture. This is advantageous because this design maintains the causality of forecasting and token-wise supervision, thereby realizing strong autoregressive capabilities, as pointed out in  [30]. To examine the benefits of the decoder-only approach, we designed an encoder-only variation of CITRAS. In this variation, we remove causal masking in cross-time attention, allowing each token to attend to others along the time dimension. Additionally, we apply a projection layer to the flattened features of all tokens in the target variable to predict future values. Table 5 and Table 6 present the comparison results between the encoder-only and decoder-only designs in covariate-informed settings (with future covariates) and multivariate settings, respectively. It is observed that in both settings, the decoder-only design outperforms the encoder-only design across most datasets. This result demonstrates the strength of CITRAS, which seamlessly integrates covariates into the decoder-only Transformer to effectively model both cross-variate and cross-time dependencies for improved forecasting accuracy.
Table 5: Comparison of encoder-only and decoder-only designs on the covariate-informed forecasting
Design
EPF-NP
EPF-PJM
EPF-BE
EPF-FR
EPF-DE
@@ B.1 Comparison between Decoder-only and Encoder-only Designs
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
Encoder-only
0.188
0.236
0.084
0.181
0.365
0.240
0.399
0.198
0.240
0.312
0.068
0.189
0.366
0.376
Decoder-only (ours)
0.171
0.214
0.063
0.153
0.362
0.248
0.363
0.176
0.217
0.293
0.070
0.193
0.280
0.313
Table 6: Comparison of encoder-only and decoder-only designs on the multivariate forecasting. The evaluation protocol is the same as in Table 2, and averaged results across the four forecasting lengths {96, 192, 336, 720} are shown.
Design
@@
Weather
PEMS08
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
Encoder-only
0.428
0.446
0.371
0.410
0.366
0.397
0.288
0.342
0.162
0.254
0.227
0.268
0.257
0.240
Decoder-only (ours)
0.389
0.421
0.359
0.405
0.350
0.382
0.262
0.320
0.160
0.253
0.233
0.272
0.236
0.216
B.2 Learnable Smoothing Factor
The smoothing factor ($\alpha$) in Attention Score Smoothing can be treated as a learnable parameter. However, as shown in Table 7, we empirically find that a fixed smoothing factor is sufficient to achieve the desired smoothing effect. Consequently, our Attention Score Smoothing introduces no new learning parameters.
Table 7: Comparison between learnable and fixed smoothing factor
Smoothing factor
EPF-NP
EDF
ETTm1
ETTm2
@@ B.2 Learnable Smoothing Factor
0.350
0.382
0.262
0.320
Appendix C Showcase
To visually demonstrate the covariate understanding capability of CITRAS, we present visualization results for the EPF-NP, EDF, and BS datasets in Figure 5. We show the forecasting results under two covariate usage settings: None (no covariates used) and P+F (past and future covariates used). It is observed that CITRAS can capture both positive and negative correlations between the target variable and covariates effectively. In the EPF-NP result, CITRAS accurately predicts the sudden rise in target electricity prices by considering the high grid load during the forecasting horizon. In the EDF result, it adjusts the electricity demand downward by recognizing that the forecasting horizon falls on a holiday. In the BS result, CITRAS effectively captures the negative correlation between working days and bike rental counts, accurately forecasting the rise in counts on the day following a working day. Interestingly, patch-based models, including CITRAS and Timer-XL, are better at reflecting the local impact of future covariates on target forecasting than TimeXer. This is likely because the variate-level approach adopted by TimeXer cannot accurately capture the local impact of covariates within the forecasting horizon, as it combines these pivotal features and past values into a single variate token.
Figure 5: Visualization results for the EPF-NP (top), EDF (middle), and BS (bottom) datasets under two covariate usage settings: None (no covariates used, colored in green) and P+F (past and future covariates used, colored in blue). Observed covariates are omitted from the visualization.
Appendix D Computational Efficiency

### v3

@@ 4.3 Model Analysis
Qualitative Analysis
To demonstrate the strong covariate understanding capability of CITRAS, we visually compare the forecasting results without and with the use of covariates, as shown in Fig. 3. In the EPF-NP result, CITRAS accurately predicts the sudden rise in target electricity prices by taking into account the high grid load and low power generation forecasts during the forecasting horizon. In the BS result, it effectively captures the negative impact of the binary working day covariate and accurately forecasts the increase in rental counts on the day following a working day. These results indicate that CITRAS does not simply incorporate covariates that resemble the target shape, but rather captures complex relationships from a global perspective and accurately reflects the local impact of future covariates. This multi-grained utilization of covariates gives CITRAS a distinct advantage over TimeXer, which leverages cross-variate dependencies only at the variate level.
Figure 3: Forecasting examples of CITRAS and TimeXer without (shown in green) and with (shown in blue) the use of covariates. The target variable is displayed in the top box, while the known covariates are shown in the boxes below.
Ablation Study
CITRAS introduces KV Shift and Attention Score Smoothing (ASS) on top of a decoder-only Transformer. To validate the effectiveness of these design choices, we conduct detailed ablation studies, as summarized in Table 3. In the KV Shift ablation, the model fails to incorporate future covariates, leading to a significant performance drop in covariate-informed forecasting. In the ASS ablation, the model is restricted to capturing cross-variate dependencies only from a local perspective, resulting in inferior performance in both covariate-informed and multivariate settings. A more in-depth analysis of the sensitivity to the smoothing factor is provided in Appendix B.1. For the decoder-only ablation, we alternatively adopt an encoder-only architecture (i.e., removing causal masking in cross-time attention and applying a projection layer to the flattened features of all tokens for prediction). However, this design cannot effectively model causality in forecasting and lacks token-wise supervision, thus failing to fully leverage the inherent autoregressive capability of the Transformer in most cases. Collectively, these results highlight the strength of CITRAS, which seamlessly integrates heterogeneous covariates into a decoder-only Transformer and effectively models both cross-variate and cross-time dependencies.
@@
(w/o) ASS
0.175
0.074
0.299
0.401
0.165
(w/o) Decoder-only
0.188
0.068
0.366
0.428
0.162
Table 3: MSE results of the ablation study.
Representation Analysis
To uncover the effects of Attention Score Smoothing further, we present the visualization of the cross-variate attention scores between the target variable and other variables in the EDF dataset in Figure 4. It is observed that patch-wise attention scores effectively capture the local negative impact of ”Holiday” on the target variable (”Electricity Demand”) . However, these scores are also susceptible to local disturbances, as demonstrated by the relatively fluctuating scores of ”Humidity” compared to the ”Temperature” that shows stable correlation with the target variable. By applying Attention Score Smoothing, the obtained scores effectively convey the strong dependency on ”Holiday” from past steps and mitigate the noisy interaction caused by the fluctuating scores of ”Humidity”. In this way, Attention Score Smoothing enjoys the advantages of both patch-level and variate-level approaches while offering improved interpretability.
Figure 4: Visualization of attention scores between the target ”Electricity Demand” and other variables observed in EDF dataset. By applying Attention Score Smoothing, the smoothened scores capture strong dependency on ”Holiday” while mitigating noisy interaction between ”Humidity”.
@@ 5 Conclusion
Considering the critical roles of covariates in practical time series forecasting, we propose CITRAS, a decoder-only Transformer that flexibly leverages multiple targets, observed covariates, and known covariates. Specifically, we introduce two novel mechanisms within the patch-wise cross-variate attention modules: Key-Value Shift, which seamlessly integrates future values of known covariates in to forecasting, and Attention Score Smoothing, which refines locally accurate dependencies into global variate-level dependencies. Experimentally, CITRAS outperforms state-of-the-art models on thirteen benchmarks from both covariate-informed and multivariate settings, demonstrating its versatile ability to leverage cross-variate and cross-time dependencies for improved forecasting accuracy. In the future, we will explore the generalization capability of CITRAS through large-scale pre-training to take advantage of its simple decoder-only architecture.
@@ A.1 Dataset Descriptions
We perform covariate-informed forecasting on seven real world datasets. They include target variable(s), observed covariates, and known covariates, thereby providing a realistic representation of practical forecasting scenarios. EPF datasets (Lago et al. 2021) consist of five subsets from different day-ahead electricity markets, each spanning six years: (1) EPF-NP documents the Nord Pool electricity market from 2013-01-01 to 2018-12-24, containing hourly electricity prices as a target variable, with corresponding grid load and wind power forecasts as known covariates. (2) EPF-PJM records the Pennsylvania-New Jersey-Maryland market from 2013-01-01 to 2018-12-24, containing zonal electricity prices in the Commonwealth Edison (COMED) as a target variable, with corresponding system load and COMED load forecasts as known covariates. (3) EPF-BE captures Belgium’s electricity market from 2011-01-09 to 2016-12-31, containing hourly electricity prices as a target variable, with corresponding load forecasts in Belgium and generation forecasts in France as known covariates. (4) EPF-FR documents the electricity market in France from 2012-01-09 to 2017-12-31, containing hourly prices as a target variable, with corresponding load and generation forecasts as known covariates. (5) EPF-DE records the German electricity market from 2012-01-09 to 2017-12-31, containing hourly prices as a target variable, with zonal load forecasts in the TSO Amprion zone and wind and solar generation forecasts as known covariates. (6) EDF dataset (Aguilar Madrid and Antonio 2021) is an hourly time series dataset spanning over five years from 2015-01-03 to 2020-06-27, where the target variable is the national electricity demand in Panama. Observed covariates include four meteorological indicators (T2M=Air temperature at 2 m ($^{\circ}\mathrm{C}$), QV2M=Specific humidity at 2 m (%), TQL=Total precipitable liquid water ($1/m^{2}$), and W2M=Wind speeed at 2 m ($m/s$)) measured in three cities (Tocumen, Santiago, David) in Panama, resulting in a total of 12 indices. In addition, we adopt three binary calendar features (holiday, school day, weekend) as known covariates. (7) BS (Fanaee-T 2013) is an hourly time series data spanning two years from 2011-01-01 to 2012-12-31, where the target variables are three types of rental counts (casual, registered, and cnt=total count) in Bike Sharing system in Washington D.C. Observed covariates include five meteorological indicators (weathersit=weather situation, temp=temperature, atemp=feeling temperature, hum=humidity, and windspeed). In addition, we adopt three calendar features (holiday, weekday, working day) as known covariates.
@@
In addition, we perform multivariate forecasting on nine real-world datasets: ETT (Zhou et al. 2021) contains 7 monitoring factors in the electricity transformers from July 2016 to July 2018. (1) ETTh1 and (2) ETTh2 are hourly subsets, and (3) ETTm1 and (4) ETTm2 are subsets recorded every 15 minutes. (5) Weather (Wu et al. 2021) includes 21 meteorological factors recorded every 10 minutes from the Weather Station of the Max Planck Biogeochemistry Institute in 2020. (6) ECL (Wu et al. 2021) comprises hourly electricity consumption data from 321 clients. (7) Traffic (Wu et al. 2021) records hourly road occupancy rates measured by 862 sensors on San Francisco Bay area freeways from January 2015 to December 2016. (8) PEMS04 (Liu et al. 2022a) contains the public traffic network data recorded at 307 locations in California every 5 minutes from 2018-01-01 to 2018-02-28. Similarly, (9) PEMS08 (Liu et al. 2022a) contains the data at 170 locations from 2016-07-01 to 2016-08-31. We use all the variables in these datasets as target variables. The summary of datasets are provided in Table 4.
@@
Covariate-informed Forecasting EPFs 1 0 2 1 Hour (36500, 5219, 10460)
@@
Multivariate Forecasting ETTh 7 0 0 1 Hour (8545, 2881, 2881)
@@
ECL 321 0 0 1 Hour (18317, 2633, 5261)
Weather 21 0 0 10 Minutes (36792, 5271, 10540)
@@ A.2 Experiment Settings
All the experiments are implemented in PyTorch (Paszke et al. 2019). We utilize a single NVIDIA A100 80GB GPU for the ECL, Traffic, and PEMS datasets and a single NVIDIA V100 32GB GPU for the other datasets. We apply Series Stationarization (Liu et al. 2022c) to the embedding and projection layers of CITRAS to mitigate nonstationarity. We adopt ADAM (Kingma and Ba 2014) with an initial learning rate of $10^{-4}$ and L2 loss for model optimization. The training process is fixed at 10 epochs with an early stopping mechanism, which terminates training if the validation performance does not improve for 3 consecutive epochs. We set the number of layers in our proposed model $L\in\{1,2,3,4\}$, the number of heads in the multi-head attention layers as 8, dimension of token embedding $D\in\{128,256,512,1024\}$, and the smoothing factor for Attention Score Smoothing $\alpha\in\{0.1,0.2,0.4,0.8\}$. The patch length is uniformly set to 24 for covariate-informed forecasting and 96 for multivariate forecasting. We ensure that the drop last option is set to False in the test data loader to avoid any unfair impact, as pointed out by Qiu et al.. All evaluations are conducted with a fixed random seed and a single run for each experiment.
@@ A.3 Baseline Descriptions
We reproduced the baseline models based on the TimesNet (Wu et al. 2023) repository or official repositories. Note that Timer-XL, TimeXer  [^1] The TimeXer implementation that utilizes future covariates is not officially released., iTransformer, and Leddam are extended from their original implementations to support future covariates in covariate-informed forecasting. Specifically, for Timer-XL, we modified the masking scheme in TimeAttention so that target token embeddings can attend to known covariate token embeddings from one patch step ahead. For TimeXer, iTransformer, and Leddam, we introduced an additional variate embedding layer for known covariates to handle their longer lengths.
@@ Appendix B Extra Model Analysis
B.1 Attention Score Smoothing
Attention Score Smoothing captures locally accurate patch-level cross-variate dependency and refines it into global variate-level dependency. Note that a large smoothing factor alpha emphasizes patch-level dependency (with $\alpha=1.0$ corresponding to no smoothing), while a smaller alpha emphasizes variate-level dependency through heavier smoothing. Figure 5 illustrates the performance at different levels of granularity in cross-variate attention. The performance of the variate-level approach is represented by TimeXer. We observe that the suitability of variate-level versus patch-level approaches is dependent on the dataset. However, strong Attention Score Smoothing (around $\alpha=0.2$) strikes a balance between the two, highlighting the versatile effectiveness of this mechanism in both covariate-informed and multivariate forecasting.
Note that this smoothing factor can be treated as a learnable parameter. However, as shown in Table 5, we empirically find that a fixed smoothing factor is sufficient to achieve the desired smoothing effect. Consequently, our Attention Score Smoothing introduces no new learning parameters.
Figure 5: The performance at different levels of granularity in cross-variate attention. The performance at ”Variate” level is represented by TimeXer. Strong Attention Score Smoothing (around $\alpha=0.2$) outperforms both variate-level and patch-level approach.
Smoothing factor
@@ B.1 Attention Score Smoothing
0.280
0.313
0.389
0.421
0.160
0.253
Table 5: Comparison between learnable and fixed smoothing factor.
B.2 Computational Efficiency